OpenStudy (anonymous):

Can someone help me on this? Given that 5a^2-4a-13-3(a^2+a+2)=0.What is the sum of the possible values of a?

OpenStudy (anonymous):

can be simplified into \[2 a^2-7 a-19\]

OpenStudy (anonymous):

When I solve, I get \[a = 1/4 (7-\sqrt{201})\] \[a = 1/4 (7+\sqrt{201})\] add them

OpenStudy (anonymous):

just as a short cut, once you have a quadratic polynomial in the form of: \[x^2+bx+c = 0\] the sum of the roots is always -b. No need to actually solve for the roots.